# Compress, Gather, and Recompute: REFORMing Long-Context Processing in Transformers

### v1

@@ 3.3 On-Demand Cache Recomputation
Identification of significant inputs. After constructing the cross-layer context embeddings corresponding to the input, we perform a token-level cosine similarity search between the query (the last part of the input) and the remaining inputs using the context embeddings. Then, we max-pool the similarity scores over the query to tokens to ensure that each token is assigned a single score. To preserve the continuity of the identified inputs, we further max-pool each token’s score with the 128 adjacent tokens. After processing the significance scores, we identify the tokens with the highest scores. We always keep the initial and final 256 tokens to maintain coherence.
@@ 4 Experiments
This section demonstrates the performance of our method across diverse tasks. In Section 4.1, we begin by showcasing the precise retrieval ability of our approach using a needle-in-a-haystack benchmark. In Section 4.2, we evaluate our method on RULER (Hsieh et al., 2024), a comprehensive benchmark comprising multiple synthetic datasets that include more diverse and challenging needle-in-a-haystack tasks, aggregation tasks, and question answering tasks. Also, we extend our evaluations to more a demanding BABILong dataset (Kuratov et al., 2024), which features complex long-context assessments, including multi-hop reasoning. In Section 4.3, we further test REFORM’s performance on more realistic tasks from $\infty$-Bench (Zhang et al., 2024). In Section 4.4, we highlight the flexibility of our approach by evaluating it on multi-modal benchmarks. In Section 4.5, we compare our approach to retrieval-augmented generation, an emerging direction for handling long inputs. Finally, we analyze the efficiency of our approach in Section 4.6.
Common setup and baselines. Throughout the paper, we mainly compare our approach against training-free methods that utilize recurrence-based and random-access approaches. Specifically, we compare our method against StreamingLLM (Xiao et al., 2023a), TOVA (Oren et al., 2024), H2O (Zhang et al., 2023), InfiniPot (Kim et al., 2024), and InfLLM (Xiao et al., 2024a). We also include a truncation baseline, which simply drops the middle part of the input. To enhance efficiency, we implement recurrent chunked forwarding for recurrence-based methods as described in Section 3.1 for all compression baselines, replacing token-level recurrence. For H2O, we restrict attention score computations to the last 128 tokens of each chunk for efficient implementation.
For all text-based experiments, we use Mistral-NeMo-Instruct-2407 (Jiang et al., 2023) and Qwen2.5-7B-Instruct (Yang et al., 2024) models. For multi-modal experiments, we use Pixtral-12B-2409 (Agrawal et al., 2024). All recurrent baselines operate with a KV cache budget and chunk size of 32k tokens, and InfLLM also uses 32k active KV cache budget. We always keep the initial and recent 256 tokens in cache for all baselines and REFORM to maintain the coherency of the text. For REFORM, we use a recomputation budget of 8k tokens for Mistral-Nemo-Instruct-2407 and 16k tokens for all other models. We provide more details in Section B.1.
@@ 4.1 Needle-In-A-Haystack Evaluation
(a) H2O
(b) Infini-Pot
(c) InfLLM
(d) REFORM (Ours)
Figure 3: Needle-In-A-Haystack Evaluation. We compare the precise retrieval ability of different long-context handling approaches by visualizing the needle-in-a-haystack performance at different depth and context lengths. All experiments were done with Qwen2.5-7B-Instruct model, and the performance is averaged over 20 samples.
To evaluate the precise retrieval performance of our approach, we employ the Needle-in-a-Haystack (NIAH) benchmark, following gkamradt (2023). In this task, a specific ”needle” sentence (“The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day”) is embedded within various depths of irrelevant context consisting of diverse essays by Paul Graham. The model must correctly answer the question: “What is the best thing to do in San Francisco?” This benchmark is more challenging than the commonly used passkey retrieval task (Mohtashami & Jaggi, 2023), where the haystack often consists of repetitive sentences, as it requires identifying and recalling a needle sentence from a highly diverse and non-repetitive context. For evaluation, we consider a response to be correct if it contains all three key phrases: “eat a sandwich”, “sit in Dolores Park”, and “a sunny day.”
In Figure 3, we compare the performance of our method against H2O, InfiniPot (representing recurrence-based methods) and InfLLM (representing random access methods). Our method demonstrates stable performance across all needle depths and context lengths up to 1M tokens, significantly outperforming both baselines. In contrast, H2O and InfiniPot exhibit forgetting issues, failing to retain critical information in longer inputs.
InfLLM also struggle at long contexts despite having the flexibility to perform random access, suggesting that it often fails to retrieve relevant information from the full KV cache. This aligns with our findings Section C.1, where certain heads indeed show poor retrieval performance. On the contrary, REFORM collects only the top-performing heads across multiple layers to construct the context embeddings, ensuring that the relevant inputs are successfully identified. These results highlight the limitations of existing approaches and underscore our method’s robustness in handling extremely long contexts while maintaining precise retrieval performance.
@@ 4.2 Performance on RULER and BABILong
Table 2: Evaluation on RULER. We measure the performance on an extended version of the RULER (Hsieh et al., 2024) benchmark, a synthetic dataset consisting of multiple needle-in-a-haystack tasks along with some aggregation and question answering tasks. We report the averaged performance of all tasks at different context lengths. The best values are highlighted in bold.
@@
64k 128k 200k 300k 400k 500k 1M Mistral-NeMo-Instruct-2407 Truncation 32.6 20.4 17.8 15.2 12.3 12.5 10.8 StreamingLLM 27.6 13.8 11.6 09.3 07.2 07.1 04.7 TOVA 21.6 15.3 14.0 11.8 07.9 08.7 04.6 H2O 15.1 07.4 07.8 05.6 04.2 05.7 03.6 InfiniPot 26.9 19.4 15.6 14.5 12.7 13.4 12.0 InfLLM 52.7 39.7 28.5 24.9 20.9 22.0 23.3 REFORM (Ours) 79.9 81.1 83.0 84.6 84.1 83.5 75.5 Qwen2.5-7B-Instruct Truncation 46.3 25.1 21.8 17.4 14.9 15.2 11.3 StreamingLLM 43.5 25.3 18.7 17.3 11.8 11.8 09.1 TOVA 66.2 27.7 25.7 25.8 21.9 20.4 17.0 H2O 51.8 20.9 18.5 17.1 11.6 12.1 08.7 InfiniPot 65.7 51.7 39.2 33.9 27.8 26.7 23.7 InfLLM 47.1 34.2 29.2 24.0 22.0 23.2 23.8 REFORM (Ours) 78.2 75.8 74.7 74.9 74.9 73.0 75.1
Table 3: Evaluation on BABILong. We measure the performance on BABILong (Kuratov et al., 2024), a synthetic benchmark consisting of more challenging tasks including multi-hop and multi-argument reasoning. We report the averaged performance of all tasks. The best values are highlighted in bold.
64k 128k 256k 512k 1M Avg. Mistral-NeMo-Instruct-2407 Truncation 32.2 26.2 17.0 13.6 14.0 20.6 StreamingLLM 38.8 23.4 15.4 11.0 06.2 19.0 TOVA 37.8 23.6 14.4 09.6 03.4 17.8 H2O 38.0 25.2 16.2 07.2 03.6 18.0 InfiniPot 39.6 26.8 18.6 11.2 08.8 21.0 InfLLM 40.6 34.0 23.6 13.0 09.6 24.2 REFORM (Ours) 57.4 51.4 50.6 47.6 48.8 51.2 Qwen2.5-7B-Instruct Truncation 48.4 33.4 27.4 20.0 15.6 29.0 StreamingLLM 53.4 40.6 33.2 23.8 19.6 34.1 TOVA 56.0 46.6 40.6 29.4 21.8 38.9 H2O 57.0 41.6 36.4 24.6 18.8 35.7 InfiniPot 59.6 51.0 53.4 48.2 40.2 50.5 InfLLM 43.0 29.2 20.4 15.4 11.4 23.9 REFORM (Ours) 61.6 60.4 59.8 58.8 58.8 59.9
@@
In this section, we further demonstrate the performance of our approach in more diverse and challenging synthetic benchmarks. Specifically, we evaluate different methods on an extended version of the RULER (Hsieh et al., 2024) and BABILong (Kuratov et al., 2024) benchmarks. RULER is a synthetic long-context benchmark consisting of multiple diverse and challenging needle-in-a-haystack tasks, as well as some aggregation and question answering tasks. BABILong further challenges the model by introducing more difficult tasks, such as multi-hop reasoning. Although the original version of RULER only supports up to 128k tokens, we further extend the dataset to 1M using the same recipe to evaluate the performance on longer inputs.
@@
We highlight the evaluation results on RULER in Table 2, and BABILong in Table 3. In both benchmarks, REFORM outperforms the baselines by a large margin, indicating its superiority in tasks that require precise recall of essential parts of the context, benefiting both from the ability to locate essential contexts from long inputs and the removal of distribution shifts in the KV cache that commonly come with recurrence-based or random-access approaches.
4.3 Performance on $\infty$-Bench
@@
Table 4: Evaluation on $\infty$-Bench. We evaluate each method on more realistic datasets from $\infty$-Bench (Zhang et al., 2024) ranging from long text summarization (En.Sum), question answering from long documents (En.QA, En.MC) and long dialogue understanding (En.Dia). The best values are highlighted in bold.
@@
En.Sum En.QA En.MC En.Dia Avg. Mistral-Nemo-Instruct-2407 Truncation 13.7 16.0 51.1 11.5 23.1 StreamingLLM 12.5 12.6 45.9 06.5 19.3 TOVA 12.3 13.8 47.2 08.0 20.3 H2O 14.2 17.6 49.3 06.0 21.8 InfiniPot 11.9 17.1 52.0 07.0 22.0 InfLLM 16.9 17.4 58.1 07.0 24.8 REFORM (Ours) 18.2 18.0 70.3 18.5 31.2 Qwen2.5-7B-Instruct Truncation 29.0 13.3 43.2 15.0 25.1 StreamingLLM 29.2 08.6 52.4 14.5 26.2 TOVA 29.4 08.6 56.8 15.0 27.4 H2O 31.0 11.0 56.3 15.5 28.5 InfiniPot 30.6 11.3 59.0 17.0 29.4 InfLLM 27.6 09.6 38.0 12.0 21.8 REFORM (Ours) 27.8 16.5 61.6 21.5 31.9
In this section, we compare the performance on more realistic long-context understanding tasks $\infty$-bench (Zhang et al., 2024), derived from books and dialogues. The average context length exceeds 100k tokens for most tasks, making it one of the longest long-context benchmarks that include realistic data. We outline the evaluation results in Table 4. REFORM shows superior performance over the baselines, demonstrating that it is also capable of handling complex problems on realistic data.
4.4 Multi-Modal Evaluation
This section highlights the broad applicability of REFORM by demonstrating its performance on multi-modal benchmarks. Operating at the transformer architecture level, REFORM can be seamlessly applied to any domain the base model supports. We evaluate Pixtral-12B-2409 (Agrawal et al., 2024) on three tasks from the MM-NIAH benchmark (Wang et al., 2024a), which examine the model’s ability to correctly locate and reason on multiple text needles hidden inside a long multi-modal document. We provide detailed experiment setup in Section B.2. The evaluation results in Table 5 demonstrate the superior performance of REFORM over the baseline approaches, highlighting the flexibility and generality of our approach.
@@ 4.4 Multi-Modal Evaluation
Table 5: Multi-modal Evaluation. We measure the performance on three datasets from MM-NIAH (Wang et al., 2024a), that assesses the model’s ability to handle text needles hidden in a long multi-modal document. We report normalized performance across input lengths to ensure equal contribution from each context length range. The best results are highlighted in bold.
Retrieval Counting Reasoning Avg. Pixtral-12B-2409 Truncate 72.2 18.7 51.2 47.4 StreamingLLM 71.9 17.8 49.8 46.5 TOVA 82.9 18.8 54.1 52.0 H2O 83.3 18.9 53.5 51.9 InfiniPot 85.4 18.8 54.7 53.0 REFORM (Ours) 89.2 22.0 61.3 57.5
@@ 4.5 Comparison to Retrieval Augmented Generation
Table 6: Comparison with RAG. We compare the performance of RAG methods and REFORM on four groups of needle-in-a-haystack datasets from RULER at 300k contexts, using Mistral-NeMo-Instruct-2407 model.
Single Multikey Multivalue Multiquery Sparse RAG 86.7 77.3 88.5 90.0 Dense RAG 87.3 57.3 82.5 78.0 REFORM 99.3 93.3 98.5 100.0 REFORM + RAG 99.3 94.7 99.0 100.0
We now compare REFORM to Retrieval Augmented Generation (RAG), a popular method for processing long inputs (Li et al., 2024b; Yu et al., 2024). RAG frameworks segment inputs into smaller chunks, which are independently encoded, and use external retrieval models to identify relevant segments. While effective in some scenarios, RAG suffers from key limitations that REFORM mitigates.
@@
In Table 6, we compare the performance of REFORM against RAG approaches using sparse and dense retrievers on the needle-in-a-haystack datasets from RULER at 300k contexts. We provide a more detailed experiment setup in Section B.3. REFORM consistently outperforms both approaches in all evaluations, demonstrating its robustness and efficiency. Furthermore, we explore a hybrid approach by combining REFORM with a dense retriever, blending REFORM ’s token-level significance scores with retrieval scores using a weighted sum (25% for the dense retriever, 75% for REFORM). This hybrid approach performs even better, highlighting the complementary strengths of REFORM and RAG.
4.6 Efficiency Analysis
In this section, we highlight the efficiency benefits of our approach. To this end, we demonstrate the compute- and memory-efficiency of our method by measuring the peak memory usage and inference time required for processing a long input. Specifically, we measure the peak memory and inference time that each method requires for generating 10 tokens conditioned on 256k tokens.

### v2

@@ 3.3 On-Demand Cache Recomputation
Identification of significant inputs. After constructing the cross-layer context embeddings corresponding to the input, we perform a token-level similarity search between the query (the last part of the input) and the remaining inputs using the context embeddings. Then, we max-pool the similarity scores over the query to tokens to ensure that each token is assigned a single score. To preserve the continuity of the identified inputs, we further max-pool each token’s score with the 128 adjacent tokens. After processing the significance scores, we identify the tokens with the highest scores. We always keep the initial and final 256 tokens to maintain coherence.
@@ 4 Experiments
This section demonstrates the performance of our method across diverse tasks. In Section˜4.1, we begin by showcasing the precise retrieval ability of our approach using a needle-in-a-haystack benchmark. In Section˜4.2, we evaluate our method on RULER (hsieh2024ruler) and BABILong (kuratov2024babilong), which are more advanced synthetic datasets consisting of more complex long-context assesments, including more challenging needle-in-a-haystack tasks, aggregation tasks, question answering, and multi-hop reasoning. In Section˜4.3, we further test REFORM’s performance on more realistic tasks including $\infty$-Bench (zhang2024bench) and RepoEval (repoeval), as well as highlighting the flexibility of our approach by evaluating it on multi-modal benchmarks. In Section˜4.4, we compare our approach to retrieval-augmented generation, an emerging direction for handling long inputs. Finally, we ablate on the key components and analyze the efficiency of our approach in Section˜4.5.
Common setup and baselines. Throughout the paper, we mainly compare our approach against context extrapolation methods that do not alter the model parameters, with a primary focus on recurrence-based and random-access approaches. Specifically, we compare our method against StreamingLLM (xiao2023efficient), TOVA (oren2024transformers), H2O (zhang2023h2o), InfiniPot (kim2024infinipot), and InfLLM (xiao2024infllm). We also include a truncation baseline, which simply drops the middle part of the input. For H2O, we restrict attention score computations to the last 128 tokens of each chunk for efficient implementation.
For all text-based experiments, we use Mistral-NeMo-Instruct-2407 (jiang2023mistral) and Qwen2.5-7B-Instruct (yang2024qwen2) models. For code completion experiments, we use Qwen2.5-Coder-1.5B and 7B models. For multi-modal experiments, we use Pixtral-12B-2409 (agrawal2024pixtral). All recurrent baselines operate with a KV cache budget and chunk size of 32k tokens, and InfLLM also uses 32k active KV cache budget. We always keep the initial and recent 256 tokens in cache for all baselines and REFORM to maintain the coherency of the text. For REFORM, we use a recomputation budget of 8k tokens for Mistral-Nemo-Instruct-2407 and 16k tokens for all other models. We provide more details in Section˜C.1.
@@ 4.1 Needle-In-A-Haystack Evaluation
Figure 3: Needle-In-A-Haystack Evaluation. We visualize the retrieval accuracy of Qwen2.5-7B-Instruct at different depth and context lengths. Performance is averaged over 20 samples.
To evaluate the precise retrieval performance of our approach, we employ the Needle-in-a-Haystack (NIAH) benchmark (gkamradt2024niah). In this task, a specific "needle" sentence (“The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day”) is embedded within various depths of irrelevant context consisting of diverse essays by Paul Graham. The model must correctly answer the question: “What is the best thing to do in San Francisco?” For evaluation, we consider a response to be correct if it contains all three key phrases: “eat a sandwich”, “sit in Dolores Park”, and “a sunny day.”
In Figure˜3, we measure the performance of our method at different context lengths and needle depths. Our method demonstrates perfect performance across all setups up to 1M tokens, highlighting our method’s robustness in handling extremely long contexts while maintaining precise retrieval performance.
@@ 4.2 Performance on RULER and BABILong
Table 2: Evaluation on RULER and BABILong. We measure the performance on an extended version of the RULER (hsieh2024ruler) and BABILong (kuratov2024babilong) benchmark. We report the averaged performance of all tasks at different context lengths. The best values are highlighted in bold.
@@
RULER BABILong 64k 128k 200k 300k 400k 500k 1M 64k 128k 256k 512k 1M Mistral-Nemo-Instruct-2407 Truncation 32.6 20.4 17.8 15.2 12.3 12.5 10.8 32.2 26.2 17.0 13.6 14.0 StreamingLLM 27.6 13.8 11.6 x9.3 x7.2 x7.1 x4.7 38.8 23.4 15.4 11.0 x6.2 TOVA 21.6 15.3 14.0 11.8 x7.9 x8.7 x4.6 37.8 23.6 14.4 x9.6 x3.4 H2O 15.1 x7.4 x7.8 x5.6 x4.2 x5.7 x3.6 38.0 25.2 16.2 x7.2 x3.6 InfiniPot 26.9 19.4 15.6 14.5 12.7 13.4 12.0 39.6 26.8 18.6 11.2 x8.8 InfLLM 52.7 39.7 28.5 24.9 20.9 22.0 23.3 40.6 34.0 23.6 13.0 x9.6 REFORM (Ours) 79.9 81.1 83.0 84.6 84.1 83.5 75.5 57.4 51.4 50.6 47.6 48.8 Qwen2.5-7B-Instruct Truncation 46.3 25.1 21.8 17.4 14.9 15.2 11.3 48.4 33.4 27.4 20.0 15.6 StreamingLLM 43.5 25.3 18.7 17.3 11.8 11.8 x9.1 53.4 40.6 33.2 23.8 19.6 TOVA 66.2 27.7 25.7 25.8 21.9 20.4 17.0 56.0 46.6 40.6 29.4 21.8 H2O 51.8 20.9 18.5 17.1 11.6 12.1 x8.7 57.0 41.6 36.4 24.6 18.8 InfiniPot 65.7 51.7 39.2 33.9 27.8 26.7 23.7 59.6 51.0 53.4 48.2 40.2 InfLLM 47.1 34.2 29.2 24.0 22.0 23.2 23.8 43.0 29.2 20.4 15.4 11.4 REFORM (Ours) 78.2 75.8 74.7 74.9 74.9 73.0 75.1 61.6 60.4 59.8 58.8 58.8
@@
In this section, we further demonstrate the performance of our approach in more diverse and challenging synthetic benchmarks. Specifically, we evaluate different methods on an extended version of the RULER (hsieh2024ruler) and BABILong (kuratov2024babilong) benchmarks. RULER is a synthetic long-context benchmark consisting of diverse and challenging needle-in-a-haystack tasks, as well as some aggregation and question answering tasks. BABILong further challenges the model by introducing more difficult tasks, such as multi-hop reasoning. Although the original version of RULER only supports up to 128k tokens, we further extend the dataset to 1M using the same recipe to evaluate on longer inputs.
@@
We highlight the evaluation results on RULER and BABILong in Table˜2. In both benchmarks, REFORM outperforms the baselines by a large margin, indicating its superiority in tasks that require precise recall of essential parts of the context, benefiting both from the ability to locate essential contexts from long inputs and the removal of distribution shifts in the KV cache that commonly come with recurrence-based or random-access approaches.
4.3 Performance on $\infty$-bench, RepoEval, and Multi-Modal Evaluations
@@
Table 3: Evaluation on $\infty$-Bench. We evaluate each method on more 10 datasets from $\infty$-Bench (zhang2024bench). We did not evaluate on C.Run and M.Calc datasets since no method was capable of achieving a nonzero score with these models. The best values are highlighted in bold.
@@
R.PK R.Num R.KV En.Sum En.QA En.MC En.Dia Zh.QA C.Debug M.Find Avg. Mistral-Nemo-Instruct-2407 Truncation 27.1 21.4 x3.6 13.7 16.0 51.1 11.5 25.2 28.9 20.3 21.9 StreamingLLM 28.1 15.3 x0.0 12.5 12.6 45.9 x6.5 19.2 27.2 x0.0 16.7 TOVA 82.2 47.0 x0.0 12.3 13.8 47.2 x8.0 x6.6 25.1 x4.6 24.7 H2O 31.5 x9.5 x0.0 14.2 17.6 49.3 x6.0 21.2 26.1 15.7 19.1 InfiniPot 84.1 13.6 x0.0 11.9 17.1 52.0 x7.0 11.3 27.4 15.7 24.0 InfLLM 100.0 100.0 x1.0 16.9 17.4 58.1 x7.0 24.5 24.1 27.1 37.6 REFORM (Ours) 100.0 100.0 88.2 18.2 18.0 70.3 18.5 26.7 25.9 36.0 50.2 Qwen2.5-7B-Instruct Truncation 27.1 27.1 x7.4 29.0 13.3 43.2 15.0 x9.34 37.1 45.7 25.4 StreamingLLM 28.8 28.8 x6.0 29.2 x8.6 52.4 14.5 x9.51 32.5 28.6 23.9 TOVA 100.0 100.0 x1.2 29.4 x8.6 56.8 15.0 10.65 34.3 42.6 39.8 H2O 93.1 85.4 x0.0 31.0 11.0 56.3 15.5 11.97 34.8 44.6 38.4 InfiniPot 100.0 99.8 x0.8 30.6 11.3 59.0 17.0 x9.99 36.6 44.9 41.0 InfLLM 100.0 99.8 x1.6 27.6 x9.6 38.0 12.0 10.41 29.7 45.1 37.4 REFORM (Ours) 100.0 100.0 32.8 27.8 16.5 61.6 21.5 11.81 33.0 21.7 42.7
@@
Table 4: Evaluation on RepoEval and $M$-NIAH. For RepoEval, we report the edit similarity (ES) score on RepoEval api-level completion task and line-level completion task with 1.5B and 7B models. For $M$-NIAH, we report normalized performance across input lengths to ensure equal contribution from each context length range. We do not run multi-modal evaluation for InfLLM, as its implementation only supports text-based models. Best results are in bold.
Method RepoEval $M$-NIAH
1.5B API 1.5B Line 7B API 7B Line Retrieval Counting Reasoning Avg.
Truncate 54.8 63.9 59.2 59.5 72.2 18.7 51.2 47.4
StreamingLLM 55.0 62.7 59.9 58.4 71.9 17.8 49.8 46.5
TOVA 54.7 62.2 59.7 59.8 82.9 18.8 54.1 52.0
H2O 55.1 63.4 61.2 59.6 83.3 18.9 53.5 51.9
InfiniPot 59.4 68.4 66.2 63.8 85.4 18.8 54.7 53.0
InfLLM 61.8 66.8 64.3 66.3 N/A N/A N/A N/A
REFORM (Ours) 65.3 72.4 68.7 69.4 89.2 22.0 61.3 57.5
In this section, we further evaluate the performance on more diverse long context handling tasks. Specifically, we evaluate the performance of REFORM on $\infty$-bench (zhang2024bench), a more realistic long-context benchmark including tasks derived from long books and dialogues, and RepoEval (repoeval), a repository-level code completion benchmark, to demonstrate that REFORM is useful in realistic tasks. Furthermore, we highlight the broad applicability of REFORM by demonstrating its performance on a multi-modal benchmark, MM-NIAH (wang2024needle).
For $\infty$-bench, we evaluate both Mistral-Nemo-Instruct-2407 and Qwen2.5-7B-Instruct models. For RepoEval, we perform evaluation using code-specific models, namely Qwen2.5-Coder-1.5B/7B-Instruct. For each sample, we provide the entire repository as the context except for the file that is being completed for the given sample. We report the edit similarity (ES) score as the evaluation metric. Finally for multi-modal evaluations, we use Pixtral-12B-2409 (agrawal2024pixtral).
As shown in Table˜3 and Table˜4, REFORM consistently outperforms all baselines in all three benchmarks. This highlights REFORM’s superior performance on realistic tasks, and its flexibility to handle diverse inputs, even across different modalities.
@@ 4.4 Comparison to Retrieval Augmented Generation
Table 5: Comparison with RAG. We compare the performance of RAG methods and REFORM on four groups of needle-in-a-haystack datasets (single, multikey, multivalue, and multiquery) from RULER at 300k contexts, using Mistral-NeMo-Instruct-2407 model.
Single M.Key M.Value M.Query Sparse RAG 86.7 77.3 88.5 90.0 Dense RAG 87.3 57.3 82.5 78.0 REFORM 99.3 93.3 98.5 100.0      + RAG 99.3 94.7 99.0 100.0
We now compare REFORM to Retrieval Augmented Generation (RAG), a popular method for processing long inputs (li2024retrieval; yu2024defense). RAG frameworks segment inputs into smaller chunks, which are independently encoded, and use external retrieval models to identify relevant segments. While effective in some scenarios, RAG suffers from key limitations.
@@
In Table˜5, we compare the performance of REFORM against RAG approaches using sparse and dense retrievers on the needle-in-a-haystack datasets from RULER at 300k contexts. We provide a more detailed experiment setup in Section˜C.4. REFORM consistently outperforms both approaches in all evaluations, demonstrating its robustness and efficiency. Furthermore, we explore a hybrid approach by combining REFORM with a dense retriever, blending REFORM ’s token-level significance scores with retrieval scores using a weighted sum (25% for the dense retriever, 75% for REFORM). This approach performs even better, highlighting the complementary strengths of REFORM and RAG.
4.5 Ablation Studies